40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Dec 4 21:33:24 2013, Koji, Update, CDS, c1scy time-over issue mitigated TST.pngCDS.pngno_glitch.png
    Reply  Thu Dec 5 12:20:09 2013, Jenne, Update, CDS, fixing c1mcs timing - go for it 
    Reply  Thu Dec 5 16:20:26 2013, Steve, Update, CDS, glitches are gone glitchesGONE.pngNOglitches.png
Message ID: 9445     Entry time: Thu Dec 5 16:20:26 2013     In reply to: 9441
Author: Steve 
Type: Update 
Category: CDS 
Subject: glitches are gone 


c1scy had frequent time-over. This caused the glitches of the OSEM damping servos.

Today Eric Q was annoyed by the glitches while he worked on the green PDH inspection at the Y-end.

In order to mitigate this issue, low priority RFM channels are moved from c1scy to c1tst.
The moved channels (see Attachment 1) are supposed to be less susceptible to the additional delay.

This modification required the following models to be modified, recompiled, reinstalled, and restarted
in the listed order:
c1als, c1sus, c1rfn, c1tst, c1scy

Now the models are are running. CDS status is all green.
The time consumption of c1scy is now ~30us (porevious ~60us)
(see Attachment 2)

I am looking at the cavity lock of TEM00 and I have witnessed no glitch any more.
In fact, the OSEM signals have no glitch. (see Attachment 3)

We still have c1mcs having regularly time-over. Can I remove the WFS->OAF connections temporarily?

 Koji cleaned up very nicely.

Attachment 1: glitchesGONE.png  31 kB  | Hide | Hide all
Attachment 2: NOglitches.png  55 kB  | Hide | Hide all
ELOG V3.1.3-